# Block triangular preconditioning for inverse source problems in time-space fractional diffusion equations

### v1

@@ 2 Notation and preliminaries
A particular type of multi-level matrix structure is found in multi-level Toeplitz matrices, whose element in position $({\boldsymbol{i}},{\boldsymbol{j}})$ depends only on the difference ${\boldsymbol{i}}-{\boldsymbol{j}}$. They can be expressed as $\begin{bmatrix}t_{{\boldsymbol{i}}-{\boldsymbol{j}}}\end{bmatrix}_{{\boldsymbol{i}},{\boldsymbol{j}}={\boldsymbol{1}}}^{{\boldsymbol{n}}}\;\in\mathbb{C}^{N({\boldsymbol{n}})\times N({\boldsymbol{n}})}$ for any fixed ${\boldsymbol{n}}\in\mathbb{N}^{d}$. In the case where $d=1$, it is simply a Toeplitz matrix with constant entries along the diagonals, while for $d=2$ it is a block Toeplitz matrix in which each block has a Toeplitz structure. More in general, a $d$-level Toeplitz matrix is a block Toeplitz matrix whose blocks have $(d-1)$-level Toeplitz structure. When the elements $t_{\boldsymbol{k}}$, ${\boldsymbol{k}}={\boldsymbol{1}}-{\boldsymbol{n}},\ldots,{\boldsymbol{n}}-{\boldsymbol{1}}$, are the Fourier coefficients of a given complex-valued function $f\in L^{1}\big{(}[-\pi,\pi]^{d}\big{)}$, the Toeplitz matrix is said to be generated by $f$.
@@ 4 Discretization
Under mild hypothesis, an error estimate is available for this formula, see [10, Lemma 2.1] or [14, Theorem 1.6.1].
@@
On the other hand, in multi-dimensional settings we compute the coefficients $a_{\boldsymbol{l}}^{(\omega)}$ numerically through the Fast Fourier Transform (FFT).
@@ 5 Matrix formulation and preconditioning
[…]
@@ 6 Numerical experiments
We conclude our analysis by numerically testing the efficiency of the proposed strategy. Our aim is to assess the preconditioner’s effectiveness in clustering the eigenvalues at 1 and improving the condition number, and also in terms of iterations and CPU time required by the GMRES method. Furthermore, we evaluate the accuracy of the method in reconstructing the source term. All tests were run on MATLAB R2022b, using a PC with the following specifications: Intel(R) Core(TM) i7-9750H CPU @2.60GHz 2.59 GHz and 16 GB of RAM.
6.1 Problem setting
@@
$\widehat{A}_{n,S}\mathbf{y}=\widehat{\mathbf{z}}$
@@
$\widehat{A}_{n,S}:=\left[\begin{array}[]{cccc}e_{0}D_{1}^{(1)}+\frac{\eta}{h^{\omega}}D_{2}^{(1)}B_{n}&\mathrm{O}_{n}&\dots&\mathrm{O}_{n}\\ e_{1}D_{1}^{(2)}&e_{0}D_{1}^{(2)}+\frac{\eta}{h^{\omega}}D_{2}^{(2)}B_{n}&\ddots&\vdots\\ \vdots&\ddots&\ddots&\mathrm{O}_{n}\\ e_{S-1}D_{1}^{(S)}&\cdots&e_{1}D_{1}^{(S)}&e_{0}D_{1}^{(S)}+\frac{\eta}{h^{\omega}}D_{2}^{(S)}B_{n}\end{array}\right],$
@@
where $\mathbf{f}$ denotes the sampling of $f(x)$ over the spatial mesh.
Then, to simulate the contamination of $\mu(x)$ by noise as in (3), we perturb it by adding a random quantity $\varepsilon\delta(x)$, where $\delta(x)$ is randomly selected in the interval $(-1,1)$ and the noise level is set to $\varepsilon=0.01\cdot\|\mu\|_{2}$. Therefore, we have $\mu_{\varepsilon}(x)=\mu(x)+\varepsilon\delta(x)$ and our regularized model problem takes the form
@@
6.2 Spectral analysis
First, let us examine the effectiveness of the preconditioner $P_{n,S}$ in reducing the condition number and clustering the eigenvalues of the coefficient matrix at 1. The regularization parameter is set to $\lambda=5\cdot 10^{-3}$.
Figure 1 shows the distribution of the eigenvalues in the complex plane for various combinations of the fractional orders $\beta$ and $\omega$, with fixed problem dimensions $n=50$ and $S=10$. These relatively small sizes were chosen to ensure that the spectral patterns are clearly visible. For larger dimensions, the spectral behavior remains qualitatively unchanged and is therefore not included. The plots align with the theoretical predictions, showing eigenvalues accumulating near 0 for $A_{n,S}$ and a tight cluster at 1 for the preconditioned matrix, with no eigenvalues close to 0 and very few outliers. Because the cluster at 1 is difficult to discern at the original scale, in the third column we included a magnified graph focused on this accumulation point.
Figure 1: Eigenvalues of $A_{n,S}$ (left column) and $P_{n,S}^{-1}A_{n,S}$ (central column) in the complex plane, for different choices of the fractional orders, with $n=50$ and $S=10$. For clarity, we also display magnified versions of the central graphs (right column), focused on the cluster at 1.
@@ 6.2 Spectral analysis
Table 1: 2-norm condition numbers of the nonpreconditioned and preconditioned coefficient matrix.
@@
$n$ $S$ $P_{n,S}$ $P_{n,S}$ $P_{n,S}$
$2^{4}$ $2^{4}$ 2021.13 62.31 731.46 85.61 1370.83 108.14
$2^{4}$ $2^{5}$ 4124.81 63.03 1499.79 86.36 3655.00 107.30
$2^{4}$ $2^{6}$ 8176.04 64.32 3083.85 87.80 9879.29 107.67
$2^{4}$ $2^{7}$ 15871.77 66.58 6351.83 90.38 26681.00 109.32
$2^{5}$ $2^{4}$ 7656.42 65.81 1511.39 97.70 1687.82 136.21
$2^{5}$ $2^{5}$ 15643.64 66.66 2787.31 98.67 4430.20 135.51
$2^{5}$ $2^{6}$ 31110.00 68.18 5206.05 100.50 11877.78 136.64
$2^{5}$ $2^{7}$ 60608.28 70.80 9855.65 103.75 31940.68 139.80
$2^{6}$ $2^{4}$ 29512.57 68.06 3762.59 107.17 2098.09 163.44
$2^{6}$ $2^{5}$ 60215.24 69.08 6383.24 108.29 5198.95 162.78
$2^{6}$ $2^{6}$ 119675.61 70.82 10876.91 110.39 13496.73 164.49
$2^{6}$ $2^{7}$ 233363.84 73.77 18732.27 114.11 35691.13 168.85
$2^{7}$ $2^{4}$ 113517.13 69.59 10325.93 114.37 2857.14 189.19
$2^{7}$ $2^{5}$ 231375.60 70.75 16832.09 115.60 6392.26 188.46
$2^{7}$ $2^{6}$ 459244.79 72.69 27133.18 117.92 15500.26 190.58
$2^{7}$ $2^{7}$ 894577.11 75.89 43743.30 122.02 39323.12 195.90
6.3 GMRES method performance
@@
Now, we solve the linear system of the form (11) arising from the discretization of the problem, evaluating the number of iterations and CPU time required by the nonpreconditioned and preconditioned GMRES method. The results are gathered in Table 2, which reports the number of iterations and CPU times (in seconds) for the nonpreconditioned and preconditioned GMRES, using the same values of $\beta$ and $\omega$ as in the previous subsection. The initial guess is set to the zero vector, the tolerance is $\texttt{tol}=10^{-8}$, and the maximum number of iterations is equal to the matrix size. The regularization parameter is set to $\lambda=5\cdot 10^{-3}$ and the noise level to $\varepsilon=0.01$.
@@
In the columns corresponding to the nonpreconditioned linear system, we observe high iteration counts, reflecting the severe ill-conditioning of the problem, and consequently we have high CPU times. Both the iteration counts and the CPU times are more sensitive to increases in $n$ than in $S$, in the sense that high values of $n$ result to more iterations and longer CPU times, even when the total matrix size is the same or even smaller due to a smaller $S$. This behavior is reasonable, considering that larger $n$ values lead to a denser coefficient matrix and, as discussed in the previous section, the computational cost is primarily driven by the spatial variable. These trends also align with Table 1, which shows that the condition number is generally more influenced by the increase in $n$ than in $S$.
@@
Turning to the columns related to the preconditioner, the iteration counts across the various choices of fractional orders align well with the spectral analysis. $P_{n,S}$ demonstrates an excellent performance, even showing optimal convergence, in the sense that the number of iterations is independent from the matrix size and remains almost constant as $n$ and $S$ increase. The corresponding CPU times remain very low, reflecting this good behavior. In accordance with the spectral analysis presented in the previous subsection, the number of iterations required by the preconditioned system is especially low when $\beta$ is close to 1 and $\omega$ is close to 2, while in the nonpreconditioned case we have particularly high iteration counts.
@@
Table 2: Iterations and CPU times required for solving the nonpreconditioned and preconditioned linear system with the GMRES method.
@@
$n$ $S$ $P_{n,S}$ $P_{n,S}$ $P_{n,S}$ $P_{n,S}$ $P_{n,S}$ $P_{n,S}$
$2^{4}$ $2^{4}$ 174 8 0.042 0.004 117 10 0.027 0.005 112 13 0.024 0.006
$2^{4}$ $2^{5}$ 257 8 0.122 0.008 146 11 0.061 0.010 154 13 0.063 0.011
$2^{4}$ $2^{6}$ 372 9 0.373 0.017 191 11 0.156 0.020 280 13 0.280 0.025
$2^{4}$ $2^{7}$ 509 9 1.187 0.040 273 11 0.524 0.049 543 13 1.420 0.056
$2^{4}$ $2^{8}$ 662 9 7.501 0.109 416 11 3.930 0.138 1024 13 17.759 0.187
$2^{5}$ $2^{4}$ 320 9 0.107 0.007 181 11 0.051 0.008 151 15 0.044 0.011
$2^{5}$ $2^{5}$ 475 8 0.398 0.012 217 11 0.113 0.015 199 15 0.101 0.020
$2^{5}$ $2^{6}$ 667 8 1.473 0.026 269 11 0.342 0.032 340 15 0.453 0.042
$2^{5}$ $2^{7}$ 928 9 12.775 0.064 338 11 1.850 0.079 658 16 5.941 0.105
$2^{5}$ $2^{8}$ 1229 9 40.527 0.170 468 12 7.664 0.204 1272 15 42.903 0.286
$2^{6}$ $2^{4}$ 605 9 0.512 0.014 281 11 0.152 0.020 193 17 0.091 0.024
$2^{6}$ $2^{5}$ 877 8 1.932 0.022 349 11 0.485 0.032 250 17 0.267 0.046
$2^{6}$ $2^{6}$ 1322 9 23.259 0.055 428 11 2.649 0.083 392 17 2.275 0.127
$2^{6}$ $2^{7}$ 1808 9 77.272 0.138 500 12 7.123 0.166 777 17 15.627 0.241
$2^{6}$ $2^{8}$ 2305 9 203.018 0.310 579 12 18.883 0.434 1475 17 98.988 0.673
$2^{7}$ $2^{4}$ 1176 9 3.419 0.079 482 12 0.589 0.095 250 18 0.190 0.148
$2^{7}$ $2^{5}$ 1782 8 44.396 0.120 581 12 5.097 0.161 307 18 1.626 0.305
$2^{7}$ $2^{6}$ 2562 9 158.544 0.316 698 12 12.857 0.426 450 18 5.729 0.561
$2^{7}$ $2^{7}$ 3381 9 425.344 0.654 821 12 31.292 1.033 865 18 32.407 1.162
$2^{7}$ $2^{8}$ 4605 9 1159.387 1.439 851 12 52.548 1.797 1658 19 167.080 2.849
$2^{8}$ $2^{4}$ 2360 9 68.175 0.174 842 12 8.799 0.221 352 19 1.810 0.343
$2^{8}$ $2^{5}$ 3482 9 271.765 0.318 982 12 24.112 0.425 401 19 4.738 0.618
$2^{8}$ $2^{6}$ 4983 9 869.231 0.611 1181 12 53.646 0.755 529 19 12.746 1.154
$2^{8}$ $2^{7}$ 6724 9 2389.703 1.366 1365 12 112.521 1.580 966 19 58.698 2.350
6.4 Accuracy of the reconstruction
Finally, we assess the accuracy of the method in reconstructing $f(x)$. Figure 2 shows the function $f(x)$ and its reconstruction under varying noise levels $\varepsilon$ and choices of regularization parameter $\lambda$, with the fractional orders fixed at $\beta=0.1$ and $\omega=1.9$.

### v2

@@ 2 Notation and preliminaries
A particular type of multi-level matrix structure is found in multi-level Toeplitz matrices, whose element in position $({\boldsymbol{i}},{\boldsymbol{j}})$ depends only on the difference ${\boldsymbol{i}}-{\boldsymbol{j}}$. They can be expressed as $\begin{bmatrix}t_{{\boldsymbol{i}}-{\boldsymbol{j}}}\end{bmatrix}_{{\boldsymbol{i}},{\boldsymbol{j}}={\boldsymbol{1}}}^{{\boldsymbol{n}}}\;\in\mathbb{C}^{N({\boldsymbol{n}})\times N({\boldsymbol{n}})}$ for any fixed ${\boldsymbol{n}}\in\mathbb{N}^{d}$. In the case where $d=1$, it is simply a Toeplitz matrix with constant entries along the diagonals, while for $d=2$ it is a block Toeplitz matrix in which each block has a Toeplitz structure. More in general, a $d$-level Toeplitz matrix is a block Toeplitz matrix whose blocks have (d1)(d-1)-level Toeplitz structure. When the elements $t_{\boldsymbol{k}}$, ${\boldsymbol{k}}={\boldsymbol{1}}-{\boldsymbol{n}},\ldots,{\boldsymbol{n}}-{\boldsymbol{1}}$, are the Fourier coefficients of a given complex-valued function $f\in L^{1}\big([-\pi,\pi]^{d}\big)$, the Toeplitz matrix is said to be generated by $f$.
@@ 4 Discretization
Under mild hypothesis, an error estimate is available for this formula, see [11, Lemma 2.1] or [15, Theorem 1.6.1].
@@
On the other hand, in multi-dimensional settings the coefficients $a_{\boldsymbol{l}}^{(\omega)}$ can be computed numerically through the Fast Fourier Transform (FFT).
@@ 5 Matrix formulation and preconditioning
where $B_{\boldsymbol{n}}$ is the $d$-level Toeplitz matrix of size $N({\boldsymbol{n}})\times N({\boldsymbol{n}})$ representing the fractional Laplacian, with elements $a_{\boldsymbol{l}}^{(\omega)}$ given by (8). In other words, $B_{\boldsymbol{n}}$ is the $d$-level Toeplitz matrix generated by the function $g(\boldsymbol{\theta})$ in (7). Note that, since $g(\boldsymbol{\theta})$ is real-valued and even, $B_{\boldsymbol{n}}$ is a real symmetric matrix.
@@ 6 Numerical experiments
We conclude our analysis with a set of numerical experiments designed to evaluate the efficiency of the proposed strategy. Our primary objective is to assess how effectively the preconditioner reduces the number of iterations required by the GMRES method. In support of this goal, we first conduct a spectral analysis of both the nonpreconditioned and preconditioned coefficient matrix, to visualize the preconditioner’s ability to cluster eigenvalues around 1 and improve the condition number. In addition, in the one dimensional case, we test the method’s accuracy in reconstructing the source term. All tests were run on MATLAB R2018b.
6.1 One dimensional setting
@@
$\widehat{A}_{{\boldsymbol{n}},S}\mathbf{y}=\widehat{\mathbf{z}}$
@@
$\widehat{A}_{{\boldsymbol{n}},S}:=\left[\begin{array}[]{cccc}e_{0}D_{1}^{(1)}+\frac{\eta}{h^{\omega}}D_{2}^{(1)}B_{\boldsymbol{n}}&\mathrm{O}_{\boldsymbol{n}}&\dots&\mathrm{O}_{\boldsymbol{n}}\\ e_{1}D_{1}^{(2)}&e_{0}D_{1}^{(2)}+\frac{\eta}{h^{\omega}}D_{2}^{(2)}B_{\boldsymbol{n}}&\ddots&\vdots\\ \vdots&\ddots&\ddots&\mathrm{O}_{\boldsymbol{n}}\\ e_{S-1}D_{1}^{(S)}&\cdots&e_{1}D_{1}^{(S)}&e_{0}D_{1}^{(S)}+\frac{\eta}{h^{\omega}}D_{2}^{(S)}B_{\boldsymbol{n}}\end{array}\right],$
@@
where $\mathbf{f}$ denotes the sampling of the source function over the spatial mesh.
Then, to simulate the contamination of $\mu(x)$ by noise as in (3), we perturb it by adding a random quantity $\varepsilon\delta(x)$, where $\delta(x)$ is randomly selected in the interval (1,1)(-1,1) and the noise level is set to $\varepsilon=0.01\cdot\|\mu\|_{2}$. Therefore, we have $\mu_{\varepsilon}(x)=\mu(x)+\varepsilon\delta(x)$ and our regularized model problem takes the form
@@
6.1.1 Spectral analysis
First, let us examine the effectiveness of the preconditioner $P_{n,S}$ in reducing the condition number and clustering the eigenvalues of the coefficient matrix at 1. In what follows, the regularization parameter is set to $\lambda=5\cdot 10^{-3}$.
Figure 1 shows the distribution of the eigenvalues in the complex plane for various combinations of the fractional orders $\beta$ and $\omega$, with fixed problem dimensions $n=2^{6}$ and $S=2^{4}$. These relatively small sizes were chosen to ensure that the spectral patterns are clearly visible. For larger dimensions, the spectral behavior remains qualitatively unchanged and is therefore not included. The plots align with the theoretical predictions, showing eigenvalues accumulating near 0 for $A_{n,S}$ and a tight cluster at 1 for the preconditioned matrix, with no eigenvalues close to 0 and few outliers. Since the cluster at 1 is difficult to discern at the original scale, in the third column we included a magnified graph focused on this accumulation point.
Figure 1: 1D case - Eigenvalues of $A_{n,S}$ (left column) and $P_{n,S}^{-1}A_{n,S}$ (central column) in the complex plane, for different choices of the fractional orders, with $n=2^{6}$ and $S=2^{4}$. For clarity, we also display magnified versions of the central graphs, focused on the cluster at 1 (right column).
@@ 6.1.1 Spectral analysis
Table 1: 1D case - 2-norm condition numbers of the nonpreconditioned and preconditioned coefficient matrix.
@@
$n$ $S$ - $P_{n,S}$ - $P_{n,S}$ - $P_{n,S}$
$2^{4}$ $2^{4}$ 2021 62 731 86 1371 108
$2^{4}$ $2^{5}$ 4125 63 1500 86 3655 107
$2^{4}$ $2^{6}$ 8176 64 3084 88 9879 108
$2^{5}$ $2^{4}$ 7656 66 1511 98 1688 136
$2^{5}$ $2^{5}$ 15644 67 2787 99 4430 136
$2^{5}$ $2^{6}$ 31110 68 5206 101 11878 137
$2^{6}$ $2^{4}$ 29513 68 3763 107 2098 163
$2^{6}$ $2^{5}$ 60215 69 6383 108 5199 163
$2^{6}$ $2^{6}$ 119676 71 10877 110 13497 164
$2^{7}$ $2^{4}$ 113517 70 10326 114 2857 189
$2^{7}$ $2^{5}$ 231376 71 16832 116 6392 188
$2^{7}$ $2^{6}$ 459245 73 27133 118 15500 191
$2^{8}$ $2^{4}$ 433429 71 29317 119 4461 212
$2^{8}$ $2^{5}$ 883159 72 47086 120 8733 211
$2^{8}$ $2^{6}$ 1751554 74 74164 123 19073 213
6.1.2 GMRES method performance
@@
Now, we solve the linear system of the form (11) arising from the discretization of the problem, evaluating the number of iterations and CPU time required by the nonpreconditioned and preconditioned GMRES method. The results are gathered in Table 2, which reports the number of iterations and CPU times (in seconds) for the nonpreconditioned and preconditioned GMRES, using the same values of $\beta$ and $\omega$ as in the previous subsection. We set the regularization parameter to $\lambda=5\cdot 10^{-3}$ and the noise level to $\varepsilon=0.01$. The initial guess is the zero vector, the tolerance is $\texttt{tol}=10^{-8}$, and the maximum number of iterations is equal to the minimum between the size of the matrix and 1000. When the maximum number of iterations is reached before convergence, in the table we write >1000>1000, without reporting the corresponding CPU time.
@@
In the columns corresponding to the nonpreconditioned linear system, we observe high iteration counts, reflecting the severe ill-conditioning of the problem, and consequently we have high CPU times. Both the iteration counts and the CPU times are more sensitive to increases in $n$ than in $S$, in the sense that high values of $n$ result in more iterations and longer CPU times, even when the total matrix size is the same or even smaller due to a smaller $S$. This behavior is reasonable, considering that larger $n$ values lead to a denser coefficient matrix and, as discussed in the previous section, the computational cost is primarily driven by the spatial variable. These trends also align with Table 1, which shows that the condition number is generally more influenced by the increase in $n$ than in $S$.
@@
Turning to the columns related to the preconditioner, the iteration counts across the various choices of fractional orders align well with the spectral analysis. $P_{n,S}$ demonstrates an excellent performance, even showing nearly optimal convergence, in the sense that the number of iterations is almost independent from the matrix size and remains practically constant as $n$ and $S$ increase. The corresponding CPU times remain very low, reflecting this good behavior. In accordance with the spectral analysis presented in the previous subsection, the number of iterations required by the preconditioned system is especially low when $\beta$ is close to 1 and $\omega$ is close to 2, while in the nonpreconditioned case we have particularly high iteration counts.
@@
Table 2: 1D case - Iterations and CPU times required for solving the nonpreconditioned and preconditioned linear system with the GMRES method.
@@
$n$ $S$ - $P_{n,S}$ - $P_{n,S}$ - $P_{n,S}$ - $P_{n,S}$ - $P_{n,S}$ - $P_{n,S}$
$2^{4}$ $2^{4}$ 175 9 0.19 0.04 117 10 0.06 0.01 112 13 0.07 0.02
$2^{4}$ $2^{5}$ 258 9 0.22 0.02 151 10 0.11 0.02 157 13 0.12 0.02
$2^{4}$ $2^{6}$ 372 8 0.69 0.03 185 11 0.25 0.04 277 13 0.43 0.05
$2^{4}$ $2^{7}$ 505 8 1.82 0.06 283 11 0.78 0.08 536 13 1.69 0.09
$2^{4}$ $2^{8}$ 657 9 11.61 0.17 413 11 4.71 0.20 >>1000 13 - 0.24
$2^{5}$ $2^{4}$ 318 8 0.21 0.01 172 11 0.08 0.02 151 15 0.07 0.02
$2^{5}$ $2^{5}$ 465 9 0.67 0.03 223 11 0.22 0.04 192 15 0.18 0.05
$2^{5}$ $2^{6}$ 690 9 2.38 0.06 264 11 0.58 0.07 339 15 0.87 0.10
$2^{5}$ $2^{7}$ 923 9 22.93 0.15 334 11 3.23 0.18 663 15 11.39 0.23
$2^{5}$ $2^{8}$ >>1000 9 - 0.37 463 12 11.95 0.46 >>1000 16 - 0.60
$2^{6}$ $2^{4}$ 600 8 0.95 0.04 288 11 0.27 0.04 192 17 0.14 0.06
$2^{6}$ $2^{5}$ 911 9 3.26 0.06 349 11 0.66 0.07 246 17 0.42 0.11
$2^{6}$ $2^{6}$ >>1000 9 - 0.14 425 11 4.10 0.16 393 17 3.44 0.23
$2^{6}$ $2^{7}$ >>1000 9 - 0.27 489 12 10.58 0.35 779 17 24.19 0.50
$2^{6}$ $2^{8}$ >>1000 9 - 0.69 575 12 24.16 0.82 >>1000 17 - 1.16
$2^{7}$ $2^{4}$ >>1000 9 - 0.12 479 12 1.07 0.16 250 18 0.38 0.22
$2^{7}$ $2^{5}$ >>1000 9 - 0.33 590 11 7.88 0.37 306 18 2.32 0.47
$2^{7}$ $2^{6}$ >>1000 9 - 0.49 710 12 19.49 0.61 454 18 8.57 0.86
$2^{7}$ $2^{7}$ >>1000 9 - 1.03 808 12 37.41 1.34 864 18 42.82 1.98
$2^{7}$ $2^{8}$ >>1000 9 - 2.28 840 12 72.49 2.88 >>1000 18 - 4.29
$2^{8}$ $2^{4}$ >>1000 9 - 0.40 839 12 17.34 0.57 351 19 3.18 0.75
$2^{8}$ $2^{5}$ >>1000 9 - 0.78 985 12 37.30 1.00 400 19 6.83 1.34
$2^{8}$ $2^{6}$ >>1000 9 - 1.40 >>1000 12 - 1.99 521 19 15.84 2.87
$2^{8}$ $2^{7}$ >>1000 9 - 2.92 >>1000 12 - 3.69 980 19 82.34 5.33
$2^{8}$ $2^{8}$ >>1000 9 - 5.99 >>1000 13 - 8.10 >>1000 19 - 11.10
6.1.3 Accuracy of the reconstruction
Finally, we assess the accuracy of the method in reconstructing $f(x)$.